OpenStudy (anonymous):

what is the soultion of the rational equation 3/x-3 = 2/x-2...is it x =5?

OpenStudy (aroub):

cross multiply, 2(x-3)=3(x-2) 2x-6=3x-6 -3x+2x=6-6 -x=0 x=0 umm..how did you get 5?

OpenStudy (anonymous):

yes, I'm also getting 0

OpenStudy (aroub):

:D